8x-y=4 and 5x+3y=17 solving by substitution.
Answer by TimothyLamb(4379)

---
8x-y=4
x+3y=17
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix-method:
x = 1.16
y = 5.28
---
system matrix
8 -1
1 3
---
inverse of system matrix
0.12 0.04
-0.04 0.32
---
determinant of system matrix = 25
---
now by substitution-method:
---
8x - y = 4
y = 8x - 4
---
x + 3y = 17
x + 3(8x - 4) = 17
x + 24x - 12 = 17
25x = 29
---
x = 29/25
x = 1.16
---
y = 8x - 4
y = 8*(29/25) - 4
---
y = 5.28
---
solution by substitution-method:
x = 1.16
y = 5.28
---
the two methods agree.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php